# Decoupling Bias, Aligning Distributions: Synergistic Fairness Optimization for Deepfake Detection

### v2

@@ 4.1 Experimental Settings
Datasets. To evaluate the fairness generalization capability of the proposed method, we trained the model on the widely used FaceForensics++ (FF++) dataset [rossler2019faceforensics++] and tested it on four benchmark datasets: FF++, DeepFake Detection Challenge (DFDC) [r10], DeepFakeDetection (DFD) [r12], and Celeb-DF [li2020celeb]. To evaluate the fairness attribute, we divided the test dataset into intra-group evaluations focusing separately on gender and race, and inter-group evaluations combining both attributes (e.g., Male-White, Female-Asian). Details of each annotated dataset are in Appendix LABEL:sec:experimental_settings.
@@ 4.3 Robustness Study
Fig. 3 presents the performance of all method under different degradation scenarios across FF++ and Celeb-df datasets, encompassing four common types of distortion: Image Compression (IC), Gaussian Noise (GN), Gaussian Blur (GB), and Block-Wise Noise (BWN). We introduce two additional metrics: $\Delta F_{FPR}$ and $\Delta{es-AUC}$. Experimental results show that our method exhibits comparable robustness to RSEF-FDD and outperforms other methods in terms of robustness. The robustness experiments conducted on the DFD and DFDC datasets are presented in the Appendix LABEL:sec:robustness_results.
@@ 4.5 Ablation Study
To thoroughly analyze the effectiveness of our proposed modules, we conduct systematic ablation studies from both structural and parametric perspectives. As illustrated in Fig. 5, we first investigate the influence of different decoupling ratios and iteration numbers within the structural fairness decoupling module on fairness performance. The results show that increasing the number of decoupling iterations initially improves the fairness performance (measured by $F_{FPR}$), but excessive decoupling leads to degradation. Empirically, the optimal trade-off is achieved when decoupling 2% of channels during the third iteration, which results in the best balance between fairness and robustness. The fairness performance of the ResNet-50 backbone is analyzed in the Appendix LABEL:sec:ablation_results.

### v3

# Fairness-Aware Deepfake Detection: Leveraging Dual-Mechanism Optimization

@@ 4.1 Experimental Settings
Datasets. To evaluate the fairness generalization capability of the proposed method, we trained the model on the widely used FaceForensics++ (FF++) dataset [rossler2019faceforensics++] and tested it on four benchmark datasets: FF++, DeepFake Detection Challenge (DFDC) [r10], DeepFakeDetection (DFD) [r12], and Celeb-DF [li2020celeb]. To evaluate the fairness attribute, we divided the test dataset into intra-group evaluations focusing separately on gender and race, and inter-group evaluations combining both attributes (e.g., Male-White, Female-Asian). Details of each annotated dataset are in Appendix A.
@@ 4.3 Robustness Study
Fig. 3 presents the performance of all method under different degradation scenarios across FF++ and Celeb-df datasets, encompassing four common types of distortion: Image Compression (IC), Gaussian Noise (GN), Gaussian Blur (GB), and Block-Wise Noise (BWN). We introduce two additional metrics: $\Delta F_{FPR}$ and $\Delta{es-AUC}$. Experimental results show that our method exhibits comparable robustness to RSEF-FDD and outperforms other methods in terms of robustness. The robustness experiments conducted on the DFD and DFDC datasets are presented in the Appendix B.
@@ 4.5 Ablation Study
To thoroughly analyze the effectiveness of our proposed modules, we conduct systematic ablation studies from both structural and parametric perspectives. As illustrated in Fig. 5, we first investigate the influence of different decoupling ratios and iteration numbers within the structural fairness decoupling module on fairness performance. The results show that increasing the number of decoupling iterations initially improves the fairness performance (measured by $F_{FPR}$), but excessive decoupling leads to degradation. Empirically, the optimal trade-off is achieved when decoupling 2% of channels during the third iteration, which results in the best balance between fairness and robustness. The fairness performance of the ResNet-50 backbone is analyzed in the Appendix C.
@@ 5 Conclusion
Appendix for “Fairness-Aware Deepfake Detection: Leveraging Dual-Mechanism Optimization”
Appendix A Additional Experimental Settings
Table A.1 reports the total numbers of training, validation, and test samples for each dataset, together with the sensitive attributes considered in our experiments. Training and validation are conducted exclusively on FF++ dataset.
Table A.1: Sample usage for training and testing on the FF++, DFD, DFDC, and Celeb-DF datasets. ‘-’ means not used.
Dataset Samples Sensitive Attributes
Train Validation Test Gender Race Intersection
FF++ 76139 25386 25401 Male, Female Asian, Black, White, Others
Male-Asian, Male-Black, Male-White, Male-Others
Female-Asian, Female-Black, Female-White, Female-Others
DFD - - 9385 Male, Female Black, White, Others
Male-Black, Male-White, Male-Others
Female-Black, Female-White, Female-Others
DFDC - - 22857 Male, Female Asian, Black, White, Others
Male-Asian, Male-Black, Male-White, Male-Others
Female-Asian, Female-Black, Female-White, Female-Others
Celeb-DF - - 28458 Male, Female Black, White, Others
Male-Black, Male-White, Male-Others
Female-Black, Female-White, Female-Others
Appendix B Additional Robustness Results
As shown in Fig. B.1, all methods are evaluated for robustness on the DFD and DFDC datasets under four types of perturbations: IC (Image Compression), GN (Gaussian Noise), GB (Gaussian Blur), and BWN (Block-wise Noise). The results show that our proposed method is more robust than the other baselines, with its performance remaining almost unchanged under most perturbations and even improving under GN perturbation on the DFD dataset.
Figure B.1: Additional robustness evaluation on the DFD and DFDC datasets. All methods are tested under four different types of perturbations.
Appendix C Additional Ablation Results
Fig. C.1 presents the analysis of how different decoupling iterations and decoupling ratios within the structural fairness decoupling module affect fairness performance. The results indicate that the optimal trade-off is achieved with a 2% decoupling ratio at the third iteration, which is consistent with the findings for the Xception backbone.
Figure C.1: Analysis of the impact of different decoupling iterations and decoupling ratios on the fairness performance ($F_{FPR}$) for ResNet-50 backnone.
Appendix D End-to-end Training Algorithm
The following presents the pseudocode of our training optimization procedure, which integrates Structural Fairness Decoupling and Global Distribution Alignment and implements them throughout the end-to-end training process.
Algorithm 1: Training Optimization
Input: Training dataset $\mathcal{D}$ with sensitive attributes, pre-trained model, max_iterations, num_epoch, num_batch, learning rate $\eta$, fairness weight $\lambda_{\text{fair}}$, decoupling ratio $pr_{c}$, Sinkhorn regularization coefficient $\varepsilon$, and a set of subgroups $\mathcal{J}$.
Output: A deepfake detection model with improved fairness, parameterized by $\theta_{l}$.
Initialization: $\theta_{0}$, $l=0$.
for $e=1$ to max_iterations do
 For each channel $k$ in the last convolutional layer, compute its fairness index $F_{k}$
  based on Eq. 2 using $\mathcal{D}$;
 Select $pr_{c}$ percent of the channels with the smallest $F_{k}$ as the decoupling index set $\mathcal{C}^{(e)}$;
 Apply channel decoupling to the last convolutional layer using $\mathcal{C}^{(e)}$;
for epoch =1=1 to num_epoch do
  for $b=1$ to num_batch do
   Sample a mini-batch $\mathcal{D}_{b}$ from $\mathcal{D}$;
   Compute the classification loss $\mathcal{L}_{\mathrm{cls}}$ on $\mathcal{D}_{b}$ based on Eq. 1;
   Sample an intersectional subgroup from $\mathcal{J}$ and obtain its predictions;
   Compute the fairness loss $\mathcal{L}_{\mathrm{fair}}$ based on Eq. 6;
   Compute the total loss $\mathcal{L}=\mathcal{L}_{\mathrm{cls}}+\lambda_{\text{fair}}\mathcal{L}_{\mathrm{fair}}$;
   Update parameters $\theta_{l+1}\leftarrow\theta_{l}-\eta\nabla_{\theta}\mathcal{L}$;
   $l\leftarrow l+1$;
  end
end
end
return $\theta_{l}$
Appendix E Experimental Results of Different Training Set
To verify that the proposed method also exhibits superior performance when trained on other datasets, we additionally train several methods on the DFDC dataset, using Xception and ResNet-50 as backbone networks, respectively. Since PG-FDD [lin2024preserving] requires a specific label for training, which is not provided in DFDC dataset, and RSEF-FDD [he2024a3] requires additional redundant samples that we are unable to construct from DFDC dataset, these two methods are not included in this comparison. The results for the Xception-based and ResNet-50-based models are reported in Tab. E.1 and Tab. E.2, respectively, and demonstrate that, when trained on DFDC dataset, our method still achieves superior fairness across the four test datasets while simultaneously attaining the highest detection performance.
Table E.1: Evaluation of methods with Xception backbone across intra-domain (DFDC) and cross-domain (FF++, Celeb-DF, DFD) datasets.
Gender Race Intersection Overall
DFDC Ori Xception 1.13 7.54 85.90 13.95 23.59 70.76 68.76 40.34 52.91 87.48
DAG-FDD WACV’24 [ju2024improving] Xception 2.27 6.69 82.27 18.43 25.47 64.28 68.90 39.55 47.44 85.31
DAW-FDD WACV’24 [ju2024improving] Xception 1.74 4.47 78.35 35.46 32.22 72.19 72.45 35.20 53.75 78.81
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 3.88 3.87 87.98 23.35 39.60 80.26 49.47 43.77 72.71 88.87
Ours Xception 0.87 3.10 88.50 7.67 22.80 82.59 17.67 27.57 72.90 89.42
FF++ Ori Xception 16.14 16.35 55.29 22.42 16.94 54.16 81.51 33.20 47.02 56.06
DAG-FDD WACV’24 [ju2024improving] Xception 18.00 23.36 56.06 17.45 11.17 50.60 69.19 29.73 46.13 56.82
DAW-FDD WACV’24 [ju2024improving] Xception 9.67 6.94 52.78 29.98 13.43 46.21 99.66 30.44 36.95 54.78
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 7.19 4.74 58.84 33.80 16.94 51.05 73.72 29.96 45.53 59.64
Ours Xception 1.01 1.33 59.82 9.00 6.26 51.34 19.60 8.34 42.82 60.51
Celeb-DF Ori Xception 9.13 15.43 62.33 11.32 16.95 57.44 14.71 16.19 57.27 66.09
DAG-FDD WACV’24 [ju2024improving] Xception 1.50 7.26 51.87 17.23 18.58 46.63 20.68 18.02 44.94 56.43
DAW-FDD WACV’24 [ju2024improving] Xception 19.77 25.13 53.36 24.58 23.85 42.21 23.83 25.55 42.93 58.82
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 13.29 13.72 49.79 29.12 17.30 46.76 15.99 17.46 40.75 54.01
Ours Xception 3.07 6.96 62.51 10.32 15.94 59.56 8.74 13.44 58.36 70.21
DFD Ori Xception 7.81 11.44 61.91 13.28 6.19 62.11 48.35 17.02 45.47 67.51
DAG-FDD WACV’24 [ju2024improving] Xception 0.36 11.35 65.12 13.02 10.39 63.57 27.30 24.24 57.27 68.20
DAW-FDD WACV’24 [ju2024improving] Xception 14.89 13.39 50.10 14.67 15.89 47.04 45.36 37.80 44.58 53.71
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 15.12 12.40 68.17 28.18 9.26 65.72 48.29 38.00 57.65 70.78
Ours Xception 6.25 10.94 68.61 12.33 6.09 66.87 19.26 12.00 60.95 70.83
Table E.2: Evaluation of methods with ResNet-50 backbone across intra-domain (DFDC) and cross-domain (FF++, Celeb-DF, DFD) datasets.
Gender Race Intersection Overall
DFDC Ori ResNet-50 3.19 12.60 91.66 10.18 28.48 83.38 39.22 40.33 69.57 92.43
DAG-FDD WACV’24 [ju2024improving] ResNet-50 2.30 14.24 91.76 9.06 27.39 81.53 27.43 38.95 66.70 92.47
DAW-FDD WACV’24 [ju2024improving] ResNet-50 2.06 12.89 90.56 11.13 28.46 80.34 29.95 39.83 64.24 92.01
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 3.88 10.87 87.98 23.35 39.60 80.26 49.47 43.77 72.71 88.87
Ours ResNet-50 1.44 10.44 92.54 7.62 28.23 84.99 20.69 37.19 72.12 93.92
FF++ Ori ResNet-50 10.77 11.89 57.26 41.89 22.92 50.34 95.46 35.10 42.76 57.66
DAG-FDD WACV’24 [ju2024improving] ResNet-50 8.68 9.44 58.72 21.09 19.06 50.24 57.90 22.92 43.89 59.22
DAW-FDD WACV’24 [ju2024improving] ResNet-50 14.85 16.54 57.38 21.41 12.57 48.82 62.30 32.43 42.63 57.61
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 8.59 4.74 58.84 33.80 16.94 51.05 73.72 29.96 45.53 59.64
Ours ResNet-50 8.49 13.02 59.89 10.52 6.05 54.18 39.70 22.29 45.75 60.25
Celeb-DF Ori ResNet-50 3.14 5.40 59.20 26.33 29.42 62.44 27.29 19.14 50.72 66.15
DAG-FDD WACV’24 [ju2024improving] ResNet-50 0.21 5.27 60.49 26.52 27.08 55.19 31.88 29.02 53.40 63.78
DAW-FDD WACV’24 [ju2024improving] ResNet-50 3.79 4.41 58.81 33.51 37.95 49.78 35.01 35.73 48.55 64.02
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 13.29 13.72 49.79 29.12 17.30 46.76 28.99 17.46 40.75 54.01
Ours ResNet-50 2.32 4.38 65.82 25.96 39.51 56.70 26.70 34.79 55.72 73.28
DFD Ori ResNet-50 11.89 7.88 66.20 11.07 24.81 53.77 73.70 43.89 46.76 63.76
DAG-FDD WACV’24 [ju2024improving] ResNet-50 5.22 6.85 57.48 6.69 17.95 50.66 34.71 26.31 46.82 58.75
DAW-FDD WACV’24 [ju2024improving] ResNet-50 18.99 18.81 60.99 4.02 22.74 52.84 50.34 45.11 49.43 64.03
Fairadapter ICASSP’25 [ding2025fairadapter] ViT-L/14 15.12 12.40 68.17 28.18 19.26 54.72 48.29 38.00 57.65 60.78
Ours ResNet-50 0.61 0.88 63.24 1.08 13.29 54.91 12.20 15.93 45.08 64.35
Appendix F Limitation and Future Work
One limitation of our method lies in its reliance on datasets with accurate demographic annotations. Currently, fairness-related annotations in existing datasets are quite limited; therefore, our fairness analysis can only focus on gender and race. In future work, we plan to enrich existing datasets with more fairness-related labels, and even construct our own dataset to enable more comprehensive and in-depth research.
Appendix G Effect of Trade-off Hyperparameter $\lambda$
To verify the effect of the trade-off hyperparameter in Eq. 7, we conducted sensitivity analysis on the FF++ dataset. Fig. G.1 shows the fairness metrics and detection metric AUC for different $\lambda$ values. The experimental results indicate that when $\lambda$ is set to 0.005, the model achieves optimal fairness performance while maintaining fair AUC scores. It is noteworthy that the analysis reveals a trade-off between fairness and AUC scores: as $\lambda$ increases from 0.003 to 0.005, AUC decreases, but fairness improves. To more clearly illustrate the relationship between each fairness metric and AUC, we present these dynamic changes separately in Fig. G.2, which displays the trend of increased AUC corresponding to reduced fairness.
Figure G.1: Sensitivity analysis of parameter $\lambda$ on the trade-off between fairness and detection accuracy for the gender attribute on FF++.
Figure G.2: Trends in Fairness Metrics vs. AUC Score. From left to right, the chart shows how fairness metrics for gender, race, and intersection attribute change with AUC, illustrating the trade-off between accuracy and fairness.
Appendix H The T-SNE Visualization of Demographic Features
We present in Fig. H.1 the t-SNE visualizations of demographic features extracted on FF++ by the vanilla Xception model and by our method, respectively. In the visualization, the different intersectional demographic groups extracted by the unconstrained Xception model form clearly separated clusters, whereas the features extracted by our method show these intersectional groups intermingled in the feature space. This indicates that our model has discovered common fingerprints across demographics, leading to a more fair representation. The t-SNE results further reveal that most subgroups in FF++ belong to the Male-White and Female-White categories, highlighting a strong dataset bias that makes fair detection particularly challenging and underscoring the necessity of our dual-mechanism synergistic optimization framework to improve fairness for minority subgroups.
Figure H.1: The T-SNE visualization of demographic features extracted from Xception and our method on FF++ dataset.